A rectangular soccer field is 20 yards longer than it is wide write an expression for the fields perimeter in terms of the width

Answer by macston(5194)

.
W=width; L=length=W+20yds
.
P=2(L+W)
P=2(W+20yds+W)
P=2(2W+20yds)
P=4W+40yds